OpenStudy (anonymous):

]Identify the correct x intercept and y intercept of the equation 2x - 3y = -18.

OpenStudy (aravindg):

write it as y=mx +c where c is y intercept

OpenStudy (anonymous):

Xint y=0, 2x-3(0)=-18 2x=-18 x= -9 Yint x=0, 2(0)-3y=-18 -3y=18 y= -6

OpenStudy (anonymous):

For x - intercept put y = 0 and find x.. For y intercept put x = 0 and find y from the equation given..

OpenStudy (aravindg):

a better way : divide entire equation by -18

OpenStudy (aravindg):

\[\huge \frac{x}{a}+\frac{y}{b}=1\] where a is x intercept b is y intercept

OpenStudy (anonymous):

x intercept = (9, 0); y intercept = (0, -6)

OpenStudy (anonymous):

Intercepts are single values and not ordered pairs.. x-intercept = -9 y-intercept = 6

OpenStudy (accessdenied):

\(\quad \boxed{ \: \textbf{Quote } \text{(waterineyes)} \\ \qquad \text{Intercepts are single values and not ordered pairs..} \quad } \) It would be correct to express \(x\)- and \(y\)- intercepts as points, since they do occur at specific points (on the \(x\)- or \(y\)- axes).